OpenStudy (anonymous):

How do I write these fractions as repeating decimals? 20/9 1/18 5/18 1/3 1/11

OpenStudy (anonymous):

type them into your calculator to find the answer: 20/9= 2.22222222 1/18= 0.05555555 5/18= 0.27777777 1/3= 0.333333333 1/11 =0.09090909

OpenStudy (anonymous):

Thank you SO much ACO111!!!!